OpenStudy (anonymous):

Convert the equation to the standard form for a parabola by completing the square on x or y as appropriate. y^2 + 2y - 2x - 3 = 0

OpenStudy (amistre64):

since we have y^2, lets assume this is asking for a y solution

OpenStudy (amistre64):

do you recall how to complete a square?

OpenStudy (anonymous):

with numbers not variables

OpenStudy (amistre64):

lets see if we can decompose this a complete square is also a perfect square; 4 is a perfect square, 9 is a perfect square, 16 is a perfect square etc... we want a poly such that: (y+n)^2 = y^2 + 2yn + n^2

OpenStudy (amistre64):

and we are given that 2yn = 2y from the setup

OpenStudy (anonymous):

okay so the n is the square

OpenStudy (amistre64):

n is the value we have to add and subtract to "complete" the square inthe original setup

OpenStudy (amistre64):

y^2 + 2y - 2x - 3 = 0 (y^2 + 2y) - 2x - 3 = 0 (y^2 + 2y +n - n ) - 2x - 3 = 0 (y^2 + 2y +n) - n - 2x - 3 = 0 (y+n)^2 - n - 2x - 3 = 0

OpenStudy (anonymous):

so n would be 3

OpenStudy (amistre64):

n=1 2yn = 2y when n=1 and i got a few typos in that last post y^2 + 2y - 2x - 3 = 0 (y^2 + 2y) - 2x - 3 = 0 (y^2 + 2y +n^2 - n^2 ) - 2x - 3 = 0 ; n^2 is the add/subtract (y^2 + 2y +n^2) - n^2 - 2x - 3 = 0 (y+n)^2 - n^2 - 2x - 3 = 0

OpenStudy (amistre64):

i forget what standard form is; thats the one where it has at most 3 terms right?

OpenStudy (anonymous):

(y+1)^2 = 2(x+2) this is an example of one of my choices

OpenStudy (amistre64):

hmm, so this might be using hte geometric definition of a parabola ...

OpenStudy (amistre64):

lets use n=1 (y+1)^2 - 1 - 2x - 3 = 0 (y+1)^2 = 1 + 2x + 3 (y+1)^2 = 2x + 4 (y+1)^2 = 2(x+2) does fit, im just unsure about the whole "standard" part :)

OpenStudy (anonymous):

okay thanks I will try that because the other choices are different and they don't make any since because they start with (y-1)^2

OpenStudy (amistre64):

as long as you posted it correctly above .... thats all i got to go by ;)

OpenStudy (anonymous):

yeah the problem is right I just didn't know if the right side was + or -

OpenStudy (anonymous):

thank you again